# Emission and Absorption Lines in Photospheric Radius Expansion Bursts of 4U 1820–30

### v1

@@ 2.3 Coadded spectra
Guichandut et al. (2021) presented the hydrodynamic simulation of super Eddington winds. They interpreted that the observed energy will close to the line energy at the rest frame where the photospheric expand to a large radius (more than 200 km). In our case, as reported by Yu et al. (2024), we found that the maximum radii of 14 bursts exceeded 100 km. However, the photospheric expends so fast during the PRE phase that we cannot obtain a significant line around the maximum radius in one burst. To address this, we stacked all burst spectra for a 0.1 s interval at the maximum radius, excluding burst # 2, to enhance the signal-to-noise ratio. A double blackbody model was applied to fit the coadded spectra. Fig. 5 shows the stacked spectrum from the strong photospheric expansion phase, with the red curve representing the best-fitting model. We found that the hotter blackbody component with a temperature $kT_{\mathrm{bb1}}=0.58\pm{0.01}$ keV, and a radius of $R_{\mathrm{bb1}}=117\pm{26}$ km, and the cooler blackbody component with a temperature of $kT_{\mathrm{bb2}}=0.17\pm{0.01}$ keV and a blackbody radius of $R_{\mathrm{bb2}}=1146\pm{335}$ km.
@@
As shown in Fig. 5, an excess near 1 keV indicates a clear emission line, while strong evidence of an absorption line near 3 keV is also observed. Additionally, a deviation near 1.6 keV suggests the need for a broad absorption line component. Significant residuals near 2.2 keV correspond to an absorption edge (2.0–2.4 keV) due to the gold M-shell reflectivity of the detector. The overall model includes four line components, achieving a fit with $\chi^{2}=288$ for 258 dof. For the emission line, the best-fit parameters are a centroid energy of $E_{1}=1.035\pm{0.011}$ keV, a width of $0.05\pm{0.01}$ keV, and a normalization of $0.32\pm{0.04}\leavevmode\nobreak\ \mathrm{photons\leavevmode\nobreak\ cm^{-2}\leavevmode\nobreak\ s^{-1}}$. For the broad absorption line, the parameters are $E_{2}=1.64\pm{0.01}$ keV, a width of $0.17\pm{0.03}$ keV, and a normalization of $0.06\pm{0.01}\leavevmode\nobreak\ \mathrm{photons\leavevmode\nobreak\ cm^{-2}\leavevmode\nobreak\ s^{-1}}$. The second absorption line at $E_{3}=2.93\pm{0.01}$ keV has a width of $0.07\pm{0.02}$ keV and a normalization of $0.065\pm{0.011}\leavevmode\nobreak\ \mathrm{photons\leavevmode\nobreak\ cm^{-2}\leavevmode\nobreak\ s^{-1}}$. In Fig. 6, we show the distributions and fitting result of the three spectral lines of the coadded spectrum. We found the significances of the emission line, $\sigma_{\rm E1}=14.2\sigma$, and the absorption lines, $\sigma_{\rm E2}=10.8\sigma$ and $\sigma_{\rm E3}=11.7\sigma$, respectively.
@@
From the time-resolved spectroscopy reported in Yu et al. (2024), a moderate expansion phase follows the strong expansion phase, with radii that are closely matched. We extracted spectra from the moderate expansion phase for a 0.2 s interval, we coadded these spectra and fitted the same continuum model as for the PRE phases. Additionally, spectra near the touchdown moment were also coadded and analyzed. Residuals from the fits for the moderate expansion phase (black, middle panel) and the touchdown phase (blue, bottom panel) are shown in Fig. 7. No line features were detected in the touchdown spectra, but a narrow emission line at $E_{\rm 1}\sim 1.024$ keV was identified in the moderate expansion phase ($R_{\rm bb}=47\pm{15}$ km), with a significance of $5.0\sigma$ and an improvement in $\chi^{2}$ of 30.
@@
Figure 5: The coadded spectroscopy of all bursts but burst # 2 at the maximum photospheric radius. The red solid line is the best-fitting model including the continuum and the emission line $E_{\rm 1}$, the absorption line $E_{\rm 2}$ and $E_{\rm 3}$, and the absorption lines at 2.2 keV (due to the reflectivity of gold M shell of the detector itself). Middle panel show the residuals of the fitting of the model without lines. The residuals of the best fit model to the data are plotted in the bottom panel.
@@
Figure 6: The distributions of the $\Delta C$ for $10^{5}$ simulations for the coadded spectra lines. From top to bottom, are the distributions of the simulations for 1 keV emission line ($E_{1}$), 1.6 keV absorption line ($E_{2}$), and 3 keV absorption line ($E_{3}$). The distributions fitted by $\chi^{2}$-distribution with two dofs (red solid lines). The arrow marks the measured $\Delta C$ from the true data.
@@
Figure 7: The coadded spectroscopy of all burst at the moderate expansion phase (black) and the touchdown moment (blue). The touchdown spectra show no evidence for line features, and the moderate expansion phase shows a narrow 1 keV emission line.
@@ 3 Discussion
In this work, we have analyzed the spectra of NICER observations between 2017 to present from 4U 1820–30, in which presented strong evidence for the existence of narrow spectral lines during the PRE phases. By using the Monte Carlo simulations Protassov et al. (2002) to estimate the significance, we identified the 1 keV line from 14 bursts and the 3 keV line from 12 burst, see Table 2. We coaded spectra of the maximum radius phase and identified three spectral lines in coaded spectra, which are the emission line at 1.034 keV, and the absorption lines at 1.56 and 2.929 keV, respectively. The most prominent of spectral features is the emission line and the significance of coadded spectrum around $9.4\sigma$. We can infer the possible origin and physically motivated based on the fitted results.
3.1 Originating from the accretion disk?
@@
The reflection origin interpretation presents several challenges. First, no emission or absorption lines were observed during burst # $\#\ 2$, as shown in Fig. 4, despite the high reflection intensity reported during this burst (Yu et al., 2024). Notably, the fraction of bolometric flux from disk reflection in this burst is consistent with those bursts where spectral lines were observed. Additionally, all spectral lines were detected exclusively during the strong PRE phase, which is significantly shorter than the duration of reflection. A second issue involves the significant and systematic spectral shifts in line energies associated with different PRE strengths. Strohmayer et al. (2019) found that the lines in stronger PRE bursts were systematically blueshifted by a factor of 1.046 compare to those observed in weaker bursts. Similarly, from the coadded spectra, we observed the 1 keV line in the maximum radius phase ($E_{1}=1.039$ keV at $R_{\rm bb}\sim 120$ km) exhibiting weaker redshifts than in the moderate expansion phase ($E_{1}=1.024$ keV at $R_{\rm bb}\sim 50$ km). Therefore, we conclude that the lines originating from the accretion disk might not fully account for the spectral features observed in our data.
@@ 3.2 Originating from the PRE winds
Substantial efforts have been made by numerous research groups over the past decades to search a spectral feature in NS surface. Cottam et al. (2002) reported the spectral lines from EXO 0748–676 during its X-ray bursts, however, the detection was quite controversial (see e.g., Galloway & Keek, 2021; Ascenzi et al., 2024). Bult et al. (2021) also detected a highly significant absorption line at 3.4 keV in a burst tail from IGR J17062–6143 observed with NICER. They claimed this line was the burning ashes on the stellar surface, such as Ca and Ti.
In UCXB 4U 1820-30, we detected a highly significant absorption line at 3 keV from 12 bursts and a narrow absorption feature at 1.6 keV in one burst. These features are plausibly associated with heavy elements in the burning ashes produced during the bursts. Possible interpretations for the 3 and 1.6 keV absorption lines include the He-like lines of S XV (2.8839, 3.0325, and 3.1013 keV) (Van Hoof, 2018), and transitions of Fe (XXIII, XXIV, and XXV), Mg XI, or Cr XXIV of $\sim 1.6$ keV, respectively (Strohmayer et al., 2019). Unlike the transient PRE-phase lines, Marino et al. (2023) found persistent features likely originate in the accretion disk, as evidenced by their stability and energy range.
@@
To summarize, it is reasonable to suggest that all detected lines originate within the photosphere. Strong PRE bursts likely drive super-Eddington winds, as hydrodynamic simulations indicate that heavy elements can be ejected into the wind, imprinting spectral lines and edges on the burst spectra (Yu & Weinberg, 2018). During the expansion phase, the photospheric radius can exceed 100 km, matching the wind solution of Guichandut et al. (2021). In our observations, the maximum radius during the PRE bursts exceeds this solution, see Fig. 8. In these burst-driven winds, line energies are influenced by gravitational redshift and Doppler blueshift (Strohmayer et al., 2019). Stronger PRE bursts, reaching larger radii, generate faster outflows and higher blueshifts, with observed blueshifts systematically around 1.046.
However, Guichandut et al. (2021) found that in steady-state radiation-driven winds, a steady-state models of radiation-driven winds, the wind velocity stabilizes at, $v_{\rm max}\sim 0.01c$. This results in shifts at the photosphere that are predominantly governed by gravitational redshift, with combined effects leading to shifts of less than $\Delta E/E\approx 0.02$. Observed energies slightly deviate from the rest frame under such conditions. For example, we detected emission lines at 1.03 keV, corresponding to 1.05 keV in the rest frame. If these emission lines have a consistent origin across different bursts, they are more likely Fe L-shell transitions with minimal gravitational redshift, see Fig. 8. Most NICER bursts align with the predicted combined shifts for Fe L-shell transitions.
@@ 3.3 The ratio of $E_{3}/E_{1}$
A 1 keV emission line was also observed in the persistent emission of IGR J17062–6143. The high-resolution spectroscopy interpreted that this feature was a superposition of several more narrow lines (Degenaar et al., 2017; Van Hoof, 2018). Bult et al. (2021) speculate that the line measured in burst is also a superposition of several more narrow lines.
Assuming that the spectral lines $E_{1}$ and $E_{3}$ were both produced by the PRE wind at the same location, the absorption and emission lines in a single spectrum would experience identical gravitational redshift and Doppler blueshift. Consequently, the observed line ratio $E_{3}/E_{1}$ would match the ratio measured in the rest frame. Using the measured values of $E_{1}$ and $E_{3}$ (see Table 2), we generated Gaussian distributions for each line based on their centroid energy and width. Monte Carlo simulations were then employed to calculate the line ratio and its uncertainty. Fig. 9 shows the resulting line ratio ($E_{3}/E_{1}$) distributions, which are well-fitted by two Gaussian functions with $\mu_{1}=2.819$ and $\sigma_{1}=0.057$ for the main peak, and $\mu_{2}=2.972$ and $\sigma_{2}=0.096$ for the second peak. If $E_{3}$ represents a single line, this suggests that the emission line ($E_{1}$) is a superposition of several narrower lines rather than a single broad feature. Based on the mean values of the two peaks, we infer an energy ratio of approximately 1.05 for two narrow lines around 1 keV.
Figure 9: The ratios of centroid energy between $E_{3}$ and $E_{1}$ at the same time. The distributions are well fitted by two Gaussian function. The main peak at 2.819, and secondary at 2.972.
@@ 3.4 The future prospective
Currently, directly and preciously identifying spectral features is challenging due to the moderate energy resolution of NICER detectors. Previous attempts to detect spectral features during X-ray bursts using XMM-Newton and Chandra, such as in EXO 0748–676 and Rapid Burster (in ’t Zand et al., 2017; Cottam et al., 2008), involved stacking numerous burst spectra. However, no spectral lines were detected. This lack of detection might be due to energy shifts of spectral lines during bursts or broadening effects caused by the rapid rotation of NSs. Several upcoming missions with high energy resolution and large collection areas, such as the Hot Universe Baryon Surveyor (HUBS) and the Advanced Telescope for High Energy Astrophysics (NewAthena; Antonelli et al., 2024; Cruise et al., 2025), aim to overcome these limitations with high energy resolution and large collection areas. HUBS, in particular, features an X-ray spectrometer capable of achieving eV-level precision in the 0.1–2.0 keV band (Cui et al., 2020a; Zhang et al., 2022; Zhao et al., 2024). It also offers a large field of view and an effective area exceeding $>500\leavevmode\nobreak\ {\rm cm^{2}}$ at 1 keV (Cui et al., 2020b). These capabilities make HUBS highly optimized for detecting emission or absorption lines.
4U 1820–30 is an excellent target for HUBS observations, given its relatively low Galactic column density and energetic X-ray bursts in the soft X-ray band. To evaluate HUBS’s ability to resolve narrow lines near 1 keV, we generated a mock spectrum using the fitted parameters of burst # 8 and the current ARF and RMF of HUBS. Using the XSPEC tool fakeit, we simulated spectra with a 1-second exposure time, incorporating two emission lines at 1.005 and 1.06 keV (both with width of 0.01 keV). The spectra were grouped with minimum counts of 20. The simulated spectrum clearly resolved the two spectral features, which were successfully fitted using a two-Gaussian model. The fitted line energies were $E_{a}=1.005\pm{0.002}\,\text{keV}$ and $E_{b}=1.061\pm{0.002}\,\text{keV}$. MC simulations estimated significances of approximately $\sim 10.4\sigma$ for $E_{a}$ and $\sim 12.3\sigma$ for $E_{b}$, see Fig. 11. By combining these measured energy shifts with the effects of rotational broadening, NS mass and radius can be accurately determined using a model-independent approach. This will provide crucial constraints on the equation of state for compact objects, offering valuable insights into their fundamental properties (see e.g., Chang et al., 2006; Jonker et al., 2008; Li et al., 2018).
@@
Figure 11: The distributions of the $\Delta C$ for $10^{5}$ simulations for the lines of the mock HUBS spectra. From top to bottom, are the distributions of the simulations for 1.005 ($E_{\rm a}$) and 1.06 keV ($E_{\rm b}$) emission lines. The distributions fitted by $\chi^{2}$-distribution with two dofs (red dashed-dotted lines). The arrow marks the measured $\Delta C$ from the true data.
During the progress of the current work, Barra et al. (2025) also analyzed the spectral lines in 12 type I X-ray bursts from the same NICER observations. Compared with their work, our data set include three more X-ray bursts from the UFA event files, see Table 1. Barra et al. (2025) applied different model to model the burst continuum spectra, and also identified these three lines. Moreover, Barra et al. (2025) obtained other emission residuals at 2 keV (Si XIV) and 2.6 keV (S XVI) using state-of-art plasma codes available within SPEX with a phenomenological continuum. Generally, our results are broadly consistent with Barra et al. (2025).

### v2

@@ 2.3 Coadded spectra
Guichandut et al. (2021) presented the hydrodynamic simulation of super Eddington winds. Their interpretation states that the observed energy will close to the line energy at the rest frame where the photospheric layer expands to a large radius (more than 200 km). In our case, as reported by Yu et al. (2024), we found that the maximum radii of 14 bursts exceeded 100 km. However, the photospheric expends so fast during the PRE phase that we cannot obtain a significant line around the maximum radius in one burst. To address this, we stacked all burst spectra for a 0.1 s interval at the maximum radius, excluding burst # 2, to enhance the signal-to-noise ratio (S/N). A double-blackbody model was applied to fit the coadded spectra. Figure 5 shows the stacked spectrum from the strong photospheric expansion phase, with the red curve representing the best-fitting model. We found that the hotter blackbody component with a temperature $kT_{\mathrm{bb1}}=0.58\pm{0.01}$ keV and a radius of $R_{\mathrm{bb1}}=117\pm{26}$ km, while the cooler blackbody component has a temperature of $kT_{\mathrm{bb2}}=0.17\pm{0.01}$ keV and a blackbody radius of $R_{\mathrm{bb2}}=1146\pm{335}$ km.
@@
As shown in Fig. 5, an excess near 1 keV indicates a clear emission line, while strong evidence of an absorption line near 3 keV is also observed. Additionally, a deviation near 1.6 keV suggests the need for a broad absorption line component. Significant residuals near 2.2 keV correspond to an absorption edge (2.0–2.4 keV) due to the gold M-shell reflectivity of the detector. The overall model includes four line components, achieving a fit with $\chi^{2}=288$ for 258 dofs. For the emission line, the best-fit parameters are a centroid energy of $E_{1}=1.035\pm{0.011}$ keV, a width of $0.05\pm{0.01}$ keV, and a normalization of $0.32\pm{0.04}\leavevmode\nobreak\ \mathrm{photons\leavevmode\nobreak\ cm^{-2}\leavevmode\nobreak\ s^{-1}}$. For the broad absorption line, the parameters are $E_{2}=1.64\pm{0.01}$ keV, a width of $0.17\pm{0.03}$ keV, and a normalization of $0.06\pm{0.01}\leavevmode\nobreak\ \mathrm{photons\leavevmode\nobreak\ cm^{-2}\leavevmode\nobreak\ s^{-1}}$. The second absorption line at $E_{3}=2.93\pm{0.01}$ keV has a width of $0.07\pm{0.02}$ keV and a normalization of $0.065\pm{0.011}\leavevmode\nobreak\ \mathrm{photons\leavevmode\nobreak\ cm^{-2}\leavevmode\nobreak\ s^{-1}}$. In Fig. 6, we show the distributions and fitting result of the three spectral lines of the coadded spectrum. We found the significances of the emission line, $\sigma_{\rm E1}=14.2\sigma$, and the absorption lines, $\sigma_{\rm E2}=10.8\sigma$ and $\sigma_{\rm E3}=11.7\sigma$, respectively.
@@
From the time-resolved spectroscopy reported in Yu et al. (2024), a moderate expansion phase follows the strong expansion phase, with radii that are closely matched. We extracted spectra from the moderate expansion phase for a 0.2 s interval, we coadded these spectra and fitted the same continuum model as for the PRE phases. Additionally, spectra near the touchdown moment were also coadded and analyzed. Residuals from the fits for the moderate expansion phase (in black, middle panel) and the touchdown phase (in blue, bottom panel) are shown in Fig. 7. No line features were detected in the touchdown spectra, but a narrow emission line at $E_{\rm 1}\sim 1.024$ keV was identified in the moderate expansion phase ($R_{\rm bb}=47\pm{15}$ km), with a significance of $5.0\sigma$ and an improvement in $\chi^{2}$ of 30.
@@
Figure 5: Coadded spectroscopy of all bursts (except burst # 2) at the maximum photospheric radius. The solid red line is the best-fitting model including the continuum and the emission line, $E_{\rm 1}$, the absorption lines, $E_{\rm 2}$ and $E_{\rm 3}$, and the absorption lines at 2.2 keV (due to the reflectivity of gold M shell of the detector itself). The middle panel shows the residuals of the fitting of the model without lines. The residuals of the best fit model to the data are plotted in the bottom panel.
@@
Figure 6: Distributions of the $\Delta C$ for $10^{5}$ simulations for the coadded spectra lines. From top to bottom: Distributions of the simulations for 1 keV emission line ($E_{1}$), 1.6 keV absorption line ($E_{2}$), and 3 keV absorption line ($E_{3}$). The distributions fitted by $\chi^{2}$-distribution with two dofs (red solid lines). The arrow marks the measured $\Delta C$ from the true data.
@@
Figure 7: Coadded spectroscopy of all burst at the moderate expansion phase (black) and the touchdown moment (blue). The touchdown spectra show no evidence of line features and the moderate expansion phase shows a narrow 1 keV emission line.
@@ 3 Discussion
In this work, we have analyzed the spectra of NICER observations between 2017 to present from 4U 1820–30, in which presented strong evidence to support the existence of narrow spectral lines during the PRE phases. By using the Monte Carlo simulations from Protassov et al. (2002) to estimate the significance, we identified the 1 keV line from 14 bursts and the 3 keV line from 12 burst (see Table 1). We coadded the spectra of the maximum radius phase and identified three spectral lines in the coadded spectra, namely: the emission line at 1.034 keV and the absorption lines at 1.56 and 2.929 keV, respectively. The most prominent of spectral features is the emission line and the significance of coadded spectrum around $9.4\sigma$. We can infer the possible origin and physically motivated based on the fit results.
3.1 Plausible origin from the accretion disk
@@
The interpretation of the reflection’s origin presents several challenges. First, no emission or absorption lines were observed during burst # $\#\ 2$ (as shown in Fig. 4), despite the high reflection intensity reported during this burst (Yu et al., 2024). Notably, the fraction of bolometric flux from disk reflection in this burst is consistent with those bursts where spectral lines were observed. Additionally, all spectral lines were detected exclusively during the strong PRE phase, which is significantly shorter than the duration of reflection. A second issue involves the significant and systematic spectral shifts in line energies associated with different PRE strengths. Strohmayer et al. (2019) found that the lines in stronger PRE bursts were systematically blueshifted by a factor of 1.046, compared to those observed in weaker bursts. Similarly, from the coadded spectra, we observed the 1 keV line in the maximum radius phase ($E_{1}=1.039$ keV at $R_{\rm bb}\sim 120$ km) exhibiting weaker redshifts than in the moderate expansion phase ($E_{1}=1.024$ keV at $R_{\rm bb}\sim 50$ km). Therefore, we conclude that the lines originating from the accretion disk might not fully account for the spectral features observed in our data.
@@ 3.2 Originating from the PRE winds
Substantial efforts have been made by numerous research groups over the past decades to search for a spectral feature in NS surface. Cottam et al. (2002) reported the spectral lines from EXO 0748–676 during its X-ray bursts; however, the detection was quite controversial (see e.g., Galloway & Keek, 2021; Ascenzi et al., 2024). Bult et al. (2021) also detected a highly significant absorption line at 3.4 keV in a burst tail from IGR J17062–6143 observed with NICER. They claimed this line was the burning ashes on the stellar surface, such as Ca and Ti.
In UCXB 4U 1820-30, we detected a highly significant absorption line at 3 keV from 12 bursts and a narrow absorption feature at 1.6 keV in one burst. These features are plausibly associated with heavy elements in the burning ashes produced during the bursts. Possible interpretations of the 3 and 1.6 keV absorption lines include the He-like lines of S XV (2.8839, 3.0325, and 3.1013 keV; Van Hoof 2018), as well as transitions of Fe (XXIII, XXIV, and XXV), Mg XI, or Cr XXIV of $\sim 1.6$ keV, respectively (Strohmayer et al., 2019). In contrast to the transient PRE-phase lines, Marino et al. (2023) found persistent features likely originate in the accretion disk, as evidenced by their stability and energy range.
@@
To summarize, it is reasonable to suggest that all detected lines originate within the photosphere. Strong PRE bursts likely drive super-Eddington winds, as hydrodynamic simulations indicate that heavy elements can be ejected into the wind, imprinting spectral lines and edges on the burst spectra (Yu & Weinberg, 2018). During the expansion phase, the photospheric radius can exceed 100 km, matching the wind solution of Guichandut et al. (2021). In our observations, the maximum radius during the PRE bursts exceeds this solution (see Fig. 8). In these burst-driven winds, line energies are influenced by gravitational redshift and Doppler blueshift (Strohmayer et al., 2019). Stronger PRE bursts, reaching larger radii, generate faster outflows and higher blueshifts, with observed blueshifts systematically around 1.046.
However, Guichandut et al. (2021) found that in steady-state radiation-driven winds within a steady-state model of radiation-driven winds, the wind velocity stabilizes at $v_{\rm max}\sim 0.01c$. This results in shifts at the photosphere that are predominantly governed by gravitational redshift, with combined effects leading to shifts of less than $\Delta E/E\approx 0.02$. Observed energies slightly deviate from the rest frame under such conditions. For example, we detected emission lines at 1.03 keV, corresponding to 1.05 keV in the rest frame. If these emission lines have a consistent origin across different bursts, they are more likely to be Fe L-shell transitions with minimal gravitational redshift (see Fig. 8). Most NICER bursts align with the predicted combined shifts for Fe L-shell transitions.
@@ 3.3 The ratio of $E_{3}/E_{1}$
A 1 keV emission line was also observed in the persistent emission of IGR J17062–6143. The high-resolution spectroscopy results led to the interpretation that that this feature was a superposition of several more narrow lines (Degenaar et al., 2017; Van Hoof, 2018). Bult et al. (2021) speculated that the line measured in burst is also a superposition of several more narrow lines.
Assuming that the spectral lines $E_{1}$ and $E_{3}$ were both produced by the PRE wind at the same location, the absorption and emission lines in a single spectrum would experience identical gravitational redshift and Doppler blueshift. Consequently, the observed line ratio $E_{3}/E_{1}$ would match the ratio measured in the rest frame. Using the measured values of $E_{1}$ and $E_{3}$ (see Table 1), we generated Gaussian distributions for each line based on their centroid energy and width. Monte Carlo simulations were then employed to calculate the line ratio and its uncertainty. Figure 9 shows the resulting line ratio ($E_{3}/E_{1}$) distributions, which are well-fitted by two Gaussian functions with $\mu_{1}=2.819$ and $\sigma_{1}=0.057$ for the main peak, and $\mu_{2}=2.972$ and $\sigma_{2}=0.096$ for the second peak. If $E_{3}$ represents a single line, this suggests that the emission line ($E_{1}$) is a superposition of several narrower lines rather than a single broad feature. Based on the mean values of the two peaks, we infer an energy ratio of approximately 1.05 for two narrow lines around 1 keV.
Figure 9: Ratios of the centroid energy between $E_{3}$ and $E_{1}$ at the same time. The distributions are well fitted by two Gaussian function. The main peak at 2.819, and secondary at 2.972.
@@ 3.4 The future prospective
At present, the direct and precious identification of spectral features is challenging due to the moderate energy resolution of NICER detectors. Previous attempts to detect spectral features during X-ray bursts using XMM-Newton and Chandra, such as in EXO 0748–676 and Rapid Burster (in ’t Zand et al., 2017; Cottam et al., 2008), involved stacking numerous burst spectra. However, no spectral lines were detected in those attempts. This lack of detection might be due to energy shifts of spectral lines during bursts or broadening effects caused by the rapid rotation of NSs. Several upcoming missions with high energy resolution and large collection areas, such as the Hot Universe Baryon Surveyor (HUBS) and the Advanced Telescope for High Energy Astrophysics (NewAthena; Antonelli et al., 2024; Cruise et al., 2025), aim to overcome these limitations with high energy resolution and large collection areas. In particular, HUBS features an X-ray spectrometer capable of achieving eV-level precision in the 0.1–2.0 keV band (Cui et al., 2020a; Zhang et al., 2022; Zhao et al., 2025). It also offers a large field of view and an effective area exceeding $>500\leavevmode\nobreak\ {\rm cm^{2}}$ at 1 keV (Cui et al., 2020b). These capabilities make HUBS highly optimized for detecting emission or absorption lines.
4U 1820–30 is an excellent target for HUBS observations, given its relatively low Galactic column density and energetic X-ray bursts in the soft X-ray band. To evaluate HUBS’s ability to resolve narrow lines near 1 keV, we generated a mock spectrum using the fitted parameters of burst #8 and the current ARF and RMF of HUBS. Using the XSPEC tool fakeit, we simulated spectra with a one-second exposure time, incorporating two emission lines at 1.005 and 1.06 keV (both with width of 0.01 keV). The spectra were grouped with minimum counts of 20. The simulated spectrum clearly resolved the two spectral features, which were successfully fitted using a two-Gaussian model. The fitted line energies were $E_{a}=1.005\pm{0.002}\,\text{keV}$ and $E_{b}=1.061\pm{0.002}\,\text{keV}$. The MC simulations estimated significances of approximately $\sim 10.4\sigma$ for $E_{a}$ and $\sim 12.3\sigma$ for $E_{b}$ (see Fig. 11). By combining these measured energy shifts with the effects of rotational broadening, the NS mass and radius can be accurately determined using a model-independent approach. This will provide crucial constraints on the equation of state for compact objects, offering valuable insights into their fundamental properties (see e.g., Chang et al., 2006; Jonker et al., 2008; Li et al., 2018).
@@
Figure 11: Distributions of the $\Delta C$ for $10^{5}$ simulations for the lines of the mock HUBS spectra. From top to bottom: Distributions of the simulations for 1.005 ($E_{\rm a}$) and 1.06 keV ($E_{\rm b}$) emission lines. The distributions fitted by $\chi^{2}$-distribution with two dofs (red dashed-dotted lines). The arrow marks the measured $\Delta C$ from the true data.
When the current work was in progress, Barra et al. (2025) was also analyzing the spectral lines in 12 type I X-ray bursts from the same NICER observations. Compared with their work, our data set include three more X-ray bursts from the UFA event files (see Table 2). Barra et al. (2025) applied different models to the burst continuum spectra and also identified these three lines. Moreover, Barra et al. (2025) obtained other emission residuals at 2 keV (Si XIV) and 2.6 keV (S XVI) using state-of-art plasma codes available within SPEX with a phenomenological continuum. Generally, our results are broadly consistent with Barra et al. (2025).